Your coin collection contains 41 1956 silver dollars. If your grandparents purchased them for their face value when they were new, how much will your collection be worth when you retire in 2042, assuming they appreciate at an annual rate of 9 percent?

  • $64,444.56

  • $65,801.29

  • $67,836.38

  • $10,924.47

  • $69,193.11
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value  
r=rate of interest  
n=time period.

A=41*(1.09)^86

=41*1654.54591

=$67836.38(Approx)